OpenStudy (anonymous):

The sum of two consecutive intergers is 59. What are the number ? A.29 and 30 B.29 and 31 C.30 and 31 D.20 and 39

OpenStudy (anonymous):

D

OpenStudy (anonymous):

20+39=59

OpenStudy (anonymous):

D

OpenStudy (danjs):

since they give you answers, you can check and see, or to do it if it is not very simple two consecutive numbers, let the first be n, the next will be n+1, consecutive means one and the next sum of consecutive integers -> n + n+1 "is" - aka equals 59 n + n + 1 = 59 2n = 58 n=58/2 = 29 that tells you n=29 and the other n+1 = 30